15. November 2007. U. Schoenwaelder; http://www.math.rwth-aachen.de/∼Ulrich.Schoenwaelder HB = Hochschulbibl. RWTH, HBZ = http://www.hbz-nrw.de/ (HBZ-CD-ROM Online), MB = Mathematikbibl., DB = Didaktikbibl. (Winter), FH = Bibl. Fachhochschule Aachen, FL = Fernleihe, IB Nr. Institutsbibliothek Nr., LB = HB–Lehrbuchsammlung, LS = HB–Lesesaal

Leonhard Euler (1707-1783)

[1] Don Albers. Buildung the Euler archive. An interview with the founders. Focus (MAA), 27(1):4–9, 2007. Dominic Klyve and Lee Stemkoski construct the Euler Archive, an online resource on the works of Leonhard Euler, currently (Jan. 07) at http://www.math.dartmouth.edu/ euler/.

[2] Don Albers. The MAA is celebrating the year of Euler. Focus (MAA), 27(1):3, 2007. The MAA is celebrating the 300th anniversary of Leonard Euler’s birth.

[3] Gerald L. Alexanderson. About the cover: Euler’s INTRODUCTIO IN ANALYSIN INFINITORUM . Bulletin AMS, 44(4):635–639, 2007. In [27].

[4] G. E. Andrews. Euler’s pentagonal number theorem. Mathematics Magazine, 56:279–284, 1983. MB: Z 167.

[5] George E. Andrews. Euler’s “De Partitio Numerorum”. Bulletin AMS, 44(4):561–573, 2007. In [27]. Chapter 16 of Euler’s Introductio in analysin infinitorum is titled ”De partitio Numerorum”. Our goal is to provide a survey of the impact of Euler’s chapter on the study of partitions in the following 250 plus years. ..

[6] E. T. Bell. Men of Mathematics. Simon and Schuster, 1965 (first published in 1937). MB:.

[7] Gerd Biegel. Leonhard Euler: 1707 - 1783; Mathematiker - Mechaniker - Physiker; Leben und Werk; [Ausstellung 15.04 - 1.07.2007, Braunschweigisches Landesmuseum]. Braunschweig: Braunschweigisches Landesmuseum, 2007. SWB.

[8] Gerd Biegel, Angela Klein, Menso Folkerts, Karin Reich, and Thomas Sonar. Euler-Ausstellung in Braunschweig. DMV-Mitteilungen, 15:60, 2007.

[9] Georg Brucker. Portrait of Leonhard Euler by Johann Georg Brucker, 1737. Bulletin AMS, 44(4):511, 2007. In [27]. Farbbild.

[10] J. J. Burckhardt, E. A. Fellmann, and W. Habicht, editors. Leonhard Euler 1707 – 1783, Beiträge zu Leben und Werk. Birkhäuser, 1983.

[11] Demetrios Christodoulou. The Euler equations of compressible fluid flow. Bulletin AMS, 44(4):581–602, 2007. In [27].

[12] Peter Constantin. On the Euler equations of incompressible fluids. Bulletin AMS, 44(4):603–621, 2007. In [27].

[13] William Dunham, editor. The Genius of Euler. Reflection on His Life and Work, Spectrum. MAA, 2007. ISBN 978-0-88385-558-4. There are 34 contributors. Celebrating Euler’s 300th birthday.

[14] William Dunham. Euler. The Master of Us All. Dolciani Mathematical Expositions. MAA, 5th print. 1999. ISBN 0-88385-328-0. Introductory biographical sketch and chapters describing his contributions to eight different topics - number theory, logarithms, infinite series, analytic number theory, complex variables, algebra, geometry, and combinatorics. Selected as an outstanding academic book by Choice Magazine. List $ 35.50.

[15] Harold M. Edwards. Euler’s definition of the derivative. Bulletin AMS, 44(4):575–580, 2007. In [27]. Euler’s method of defining the derivative of a function .. calls for rewriting the difference quotient in a way that remains meaningful when the denominator is zero.

[16] L. Euler. Vollständige Anleitung zur Algebra mit den Zusätzen von Joseph Louis Lagrange. Ser. 1. Opera mathematica Vol. 1. Teubner, 1911. HB: BB21-1,1. Vgl. [18].

[17] L. Euler. Briefe an eine deutsche Prinzessin über verschiedene Gegenstände aus der Physik und Philosophie. Vieweg, 1984. Philos.B.

[18] L. Euler. Elements of Algebra. Springer-Verlag, 1984. ISBN 0-387-96014-7. MR: This is a facsimile reprint of John Hewlett’s 1840 translation of Euler’s Algebra and Lagrange’s Addition thereto.. It is accompanied by an excerpt of Horner’s memoir on the life of Euler, and a eulogy by Truesdell, with a useful bibliography. Reprint of the 5th edition Longman, Orme, and Co., London 1840. Vgl. [16].

[19] L. Euler. Introduction to Analysis of the Infinite, Books I and II (J. D. Blanton, transl.). Springer–Verlag, I: 1988, II: 1990. ISBN I: 3-540-96824-5 (Berlin ...) Pp., 0-387-96824-5 (New York ...) Pp. II: 0-387-97132-7. MB: 14621. HBZ.

[20] Leonhard Euler. Anleitung zur Differentialrechnung (1790). In J. A. C. Michelsen, editor, Leonhard Eulers vollständige Anleitung zur Differenzial-Rechnung aus d. Latein., übers. u. mit Anm. u. Zusätzen begleitet von Johann Andreas Christian Michelsen. Lagarde u. Friedrich. HBZ: 361-BI. Siehe auch [21].

[21] Leonhard Euler. Leonhard Euler’s Vollständige Anleitung zur Differenzial-Rechnung / aus d. Lat. bers. u. mit Anm. u. Zusätzen begleitet von Johann Andreas Christian Michelsen / Neudruck der Ausgabe Berlin 1798; Suppl. verfasst von Johann Philipp Grueson. Band 1 - 4. Wiesbaden: LTR-Verlag, 1981. ISBN 3-88706-047-4 Pp. HBZ. Siehe auch [20].

[22] Leonhard Euler. Einleitung in die Analysis des Unendlichen. Teil 1. Berling: Springer-Verlag, 1983. ISBN 3-540-12218-4. Reprint of the 1885 edition. With an introduction by Wolfgang Walter. Review (C. J. Scriba): Bulletin AMS 44:4 (2007), 643–644; from MathSciNet, June 2007 (in [27]).

[23] Leonhard Euler. Introduction à l’analyse infinitésimale I, II. Paris: A.C.L. Ed., 1987. I: 2-87694-004-3, II: 2-87694-007-8. Nachdr. d. Ausg. Paris 1796: Übers. von Ïntroductio in analysin infinitorum I, II (1748)”durch J. B. Labey. HBZ.

[24] Emil A. Fellmann. Leonhard Euler. Rohwohlts Monographien 387. Rowohlt, 1995. ISBN 3-499-50287-5. HBZ. Vgl. [25].

[25] Emil A. Fellmann. Leonhard Euler. Birkhäuser, 2007. ISBN 978-3-7643-7538-6. HBZ. Vgl. [24].

[26] Susan Friedlander. Commentary. Bulletin AMS, 44(4):513–514, 2007. Commentary on this special Euler issue [27].

[27] Susan Friedlander, editor. Leonhard Euler Issue, Bulletin AMS 44:4. AMS, 2007. Contains articles on Leonhard Euler: [9, 26, 43, 40, 5, 15, 11, 12, 3].

[28] Fernando Q. Gouvêa. Euler’s convincing non-proofs. Focus (MAA), 27(1):10–11, 2007. On April 15, 2007, the mathematical world will celebrate the 300th anniversary of the birth of Leonhard Euler.

[29] Fernando Q. Gouvêa. Leonhard Euler at 300. Focus (MAA), 27(1):3, 2007. On April 15, 2007, the mathematical world will celebrate the 300th anniversary of the birth of Leonhard Euler.

[30] Julian Havil. Gamma: Eulers Konstante, Primzahlstrände und die Riemannsche Vermutung; mit 24 Tabellen. Berlin [u.a.]: Springer, 2007. ISBN 978-3-540-48495-0. HBZ. Euler-Konstante.

[31] A. Heyne, A. K. Heyne, and E. Pini. Leonhard Euler - Ein Mann, mit dem man rechnen kann. Birkhäuser, 2007. ISBN 978-3-7643-7779-3. HBZ. Leonhard Euler, a Man to Be Reckoned With; Birkhäuser, ISBN 978-3-7643-8332-9.

[32] Marcel Jenni. Leonhard Euler: 707-1783; Beitr. zu Leben u. Werk: Gedenkband d. Kantons Basel-Stadt. Basel: Birkhäuser, 1983. HB: BB1567.

[33] Franz Lemmermeyer. Reciprocity Laws, from Euler to Eisenstein. Springer monographs in mathematics. Berlin: Springer-Verlag, 2000. ISBN 3-540-66957-4. MB: AL 583. HB: BB5098. Review (Imamoglu): Bulletin AMS 44:4 (2007), 647–652 (in [27]).

[34] L. Mejlbo. Historical thoughts on infinite numbers. In F. Swetz, J. Fauvel, O. Bekken, B. Johansson, and V. Katz, editors, Learn from the Masters!, Classroom resource materials, chapter 14, pages 181–187. Math. Association of America, 1995. ISBN 0-88385-703-0. HB: Bb2023. Euler rechnet mit unendlichen (natürlichen) Zahlen. Die moderne Interpretation ist nicht ausgeführt.

[35] A. Pringsheim. Euler und das ‘Cauchy’sche’ Konvergenskriterium. Bibliotheca mathematica, 3(6):252–256, 1905.

[36] Edward Sandifer. The Early Mathematics of Leonhard Euler. Spectrum. The MAA tercentenary Euler celebration. MAA, 2007. ISBN 978-0-88385-559-1. HBZ. The book is a protrait of the world’s most exciting mathematics between 1725 and 1741.

[37] Man-Keung Siu. Euler and heuristic reasoning. In F. Swetz et al., editor, Learn from the Masters!, Classroom Resource Materials, pages 145–160. MAA, 1995. HB: Bb 2023.

[38] Rüdiger Thiele. Leonhard Euler 15. April 1707 - 18. September 1783. Zur Erinnerung an seinen 300. Geburtstag. DMV-Mitteilungen, 15(2):93–103, 2007.

[39] W. Tietze. Wie Leonhard Euler auf die Reihe für den Tangens geführt wurde. Der mathematische und naturwissenschaftliche Unterricht (MNU), 19:365–368, 1966/67. HB: Z848-19. Potenzsummen, Bernoulli-Zahlen.

[40] Burt Totaro. Euler and algebraic geometry. Bulletin AMS, 44(4):541–559, 2007. In [27].

[41] C. Truesdell. Leonhard Euler, supreme geometer. In C. Truesdell, editor, An Idiot’s Fugitive Essay’s on Science, pages 337–379. Springer–Verlag, 1984.

[42] V. S. Varadarajan. Euler Through Time. An New Look at Old Themes. AMS, 2006. ISBN 978-0-8218-3580-7. Euler products, Langlands program.

[43] V. S. Varadarajan. Euler and his work on infinite series. Bulletin AMS, 44(4):515–539, 2007. In [27].

[44] A. P. Youschkevitch. Euler. In Dictionary of Scientific Biography, Vol. 4. Scribner’s, New York, 1970–1980.